Suppose that you are in a 10-m-long hallway with all doors closed. Model the hallway as a tube closed at both ends. The speed of sound in air is 343 m/s.

Determine the fundamental frequency of the hallway.

Determine the first overtone frequency of the hallway.
Answer #1

Ans:-

Fundamental frequency = n = v/λ= v/(2l) = 343/20 = 17.15 Hz.

Frequency of first overtone = 2n = 2*17.15 = 34.3Hz.